4. if sin θ and cos θ are the roots of the equations ax2-bx2+c = 0, then a, b and c satisfy the relation. (a) a 2 +b 2 +2ac=0 (b) a 2 -b 2 +2ac=0 (c) a 2 +c 2 +2ac=0 (d) a 2 -b 2 -2ac=0

4.     if sin θ and cos θ are the roots of the equations ax2-bx2+c = 0, then a, b and c satisfy the relation.


(a)    a2+b2+2ac=0


(b)    a2-b2+2ac=0


(c)    a2+c2+2ac=0


(d)    a2-b2-2ac=0

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

sin2(theta) + cos2(theta) = (sinθ +cosθ)2 -2sin(theta)cos(theta) =1

sin(theta) +cos(theta) = -b/a, sin(theta)cos(theta) =c/a

therefore b2/a2 -2c/a =1, a2-b2 +2ac=0 

'Win exciting gifts by answering the questions on Discussion Forum.'